OpenStudy (anonymous):

limx→∞−14x/2+2x= limx→−∞9x−7/x^3+2x−9= limx→∞x^2−15x−8/14−8x^2= limx→∞sqrtx^2+10x/√13−3x= limx→−∞sqrtx^2+10x/√13−3x= finding the limit for each of these

OpenStudy (mandre):

\[limit_{x \rightarrow \infty}\frac{ -14x }{ 2+2x }\] Is that the first one?

OpenStudy (mandre):

\[limit _{x \rightarrow \infty}[(\frac{ -14x }{ 2+2x }).(\frac{ 1/x }{ 1/x })]\] You basically divide each term by the highest factor of x found in the denominator. \[limit _{x \rightarrow \infty}\frac{ constant }{ x ^{n} } = 0\]

OpenStudy (anonymous):

i need the answers for the last two